# The electrostatic charge on a falling water drop

### v1

@@ III Results and Discussion
Generally, the data for both syringe tips can be fit to a linear relationship. The intercept with the vertical axis ($V_{\text{tip}}=$ 0) agrees well with our measurements with a grounded tip (Fig. 2)–that is, it represents the Volta potential, $\Delta\psi$. Interestingly, the slope of the linear fit is not unity: a small increase in $V_{\text{tip}}$ leads to a slightly smaller increase in $V_{D}$. This result may be expected since the applied voltage can change the distribution of electrons at the surface and the water dipole contribution to the Volta potential in Eq. 2 Mohandas et al. (2024). Although we did not perform independent measurements of $\Delta\psi$ using alternative experimental techniques Le et al. (2017), the consistency of $V_{D}$ over multiple experimental conditions and its quantitative agreement with the expected values of $\Delta\psi$ strongly suggest their equivalency.
@@
Figure 5: PDFs of $\Delta q/\Delta m$ for different fluids dispensed from a 5 cm long, 1.60 or 2.39 mm ID stainless steel luer lock syringe tip. The salt solution and glycerol data were acquired at a flow rate of 350 $\mu$L/min on the 2.39 mm ID tip, while the high density and much lower surface tension of Fluorinert FC-70 required a flow rate of 35 $\mu$L/min on the 1.60 mm ID tip. Each distribution is computed over $n>$ 360 data points. (a) For salt water, $\langle\Delta q\rangle=$0.08 $\pm$ 0.02 pC and $\langle\Delta m\rangle=$ 0.054 $\pm$ 0.004 g. Data were captured at 48% RH. (b) For glycerol, $\langle\Delta q\rangle=$0.13 $\pm$ 0.02 pC and $\langle\Delta m\rangle=$ 0.049 $\pm$ 0.002 g. Data were captured at 48% RH. (c) For Fluorinert FC-70, $\langle\Delta q\rangle=$0.17 $\pm$ 0.03 pC and $\langle\Delta m\rangle=$ 0.02 $\pm$ 0.02 g. Data were captured at 29% RH. Overall, the addition of ions (salt) in the fluid narrows the distribution, whereas non-polar and highly insulating fluids display broader distributions.
@@
More broadly, the large potentials acquired by fluids with low conductivity evince the persistent electrostatic hazards present across a number of fields. Beyond industrial settings, however, the frictional charging of dielectric fluids may have important implications for geophysical processes on worlds with exotic potamologies. Saturn’s moon Titan, for instance, hosts extensive river and lake systems made not of water, but of liquid methane and ethane Mitri et al. (2007). While previous work has suggested that the charging of hydrocarbon solids can impact the transport of dust on Titan, whether or not flow electrification impacts the transport of hydrocarbon liquids on Titan remains unexplored.
@@
Instruments for handling of liquids (e.g. pipette tips and syringe bodies) commonly employ plastic components. In solid-solid interactions, plastics can often facilitate the transfer of large amounts of electrostatic charge. Beyond our experiments with glass syringes and metal needles, we also investigated the charge gained by water dispensed from a polypropylene (PPL) syringe body with a 5 cm long stainless steel tip, and water deposited from a glass syringe with a PTFE tip. For these experiments, we used syringes sourced from their original sterile packaging without additional cleaning. Figure 6a shows that drops dispensed consecutively from a glass syringe/metal tip combination over a period of $\approx$ 45 minutes have $\langle\Delta q/\Delta m\rangle=$2.3 $\pm$ 0.3 pC/g. These time series are consistent with the data presented in Fig. 2. In contrast, drops sourced from a PPL syringe and metal tip initially gained positive charge, with $\Delta q/\Delta m=$ 10–30 pC/g (Fig. 6b). Moreover, we observed an evolution of $\Delta q/\Delta m$ over long timescales. During some trials, $\Delta q/\Delta m$ increased by up to a factor of 2 over the length of the experiment, whereas in other cases we observed a small decrease in $\Delta q/\Delta m$. These experiments show that despite the water passing through 5 cm of metal prior to falling through air, the initial contact with the plastic syringe body dominated the sign, magnitude, and time evolution of the drop charge.
Figure 6: Time series of $\Delta q/\Delta m$ for different syringe and tip components using 15 mL of DI water at a flow rate of 350 $\mu$L/min. The expected $\Delta q/\Delta m$ range of -5 pC/g to -1 pC/g for metal and glass is highlighted by the translucent red region on each plot. Vertical axis limits vary between plots to better depict the magnitude of $\Delta q/\Delta m$. (a) Plot of $\Delta q/\Delta m$ versus time for drops dispensed from glass syringe with different metallic syringe tips. Experiments were performed between 30 and 48% RH. The tip length varied, yet all tips had a 2.69 mm ID. (b) Plot of $\Delta q/\Delta m$ versus time for drops dispensed from stainless steel tips mounted on PPL syringe bodies. All tips had a length of 5 cm and 1.70 mm ID, and experiments were performed at 46% RH. (c) Plot of $\Delta q/\Delta m$ versus time for drops dispensed from PTFE tips mounted on a glass syringe body. All tips were cut to a length of 2 cm before Neutrad cleaning and had a 1.35 mm ID, and experiments were performed at $<$ 30% RH. Overall, the introduction of plastic components reveals time and history dependent charging behavior.
@@
The fact that water drops dispensed from plastic reservoirs or syringe tips gain large amounts of charge is consistent with previous experiments. For example, water drops deposited from a pipette tip can display a large positive charge, $\sim$100 pC or more Choi et al. (2013). A more surprising result is the fact that we observe bipolar charging on water droplets interacting with PTFE tips. Like many plastics, PTFE is on the extreme lower end of the triboelectric series and is often considered to be one of the substances that most effectively gains negative charge during frictional interactions. Indeed, the positive electrification of droplets flowing on PTFE surfaces has served as the basis for a number of proposed triboelectric nano generators Kwak et al. (2016); Wu et al. (2021).
In general, our data does not provide a satisfactory explanation for the evolution of $\Delta q/\Delta m$ when a plastic component is introduced into the system. For solid-solid contacts involving hydrocarbons, microscale chemical heterogeneity along solid surfaces can lead to a large variability in charging behavior Grosjean and Waitukaitis (2023b); Sobolev et al. (2022). Moreover, history dependence in solid-solid tribocharging has recently been demonstrated during repeated contacts between a sphere and a planar surface using acoustic levitation Grosjean and Waitukaitis (2023a). In those experiments, hysteresis in water absorption was suspected to be the primary cause of history dependence. Surprisingly, materials discharged and retested under the same humidity conditions could produce different magnitudes and signs of charge transfer. Similar variability could operate in solid-liquid contacts, explaining the apparent randomness of charge gain we observed in our experiments with plastics. Lastly, surfactants leached from plastic surfaces could also contribute to the diversity of charging behaviors reported here. We suspect that the matter of water-plastic electrification will find a more satisfactory answer in future experiments considering a broader ensemble of plastics.
@@ Conclusions
Tables such as the “triboelectric series” can sort the charging behavior of different materials empirically. However, these approaches provides little physical insight into the microphysical processes underpinning contact and flow electrification. Ultimately, the charge exchanged by materials in contact may reflect a superposition of a multitude of electrification mechanisms. By quasistatically depositing individual drops at low flow rates, we show that, for certain material combinations, this charging can be attributed to well-characterized electrochemical processes. Specifically, the charge on droplets falling from a glass-metal vessel can be described by the Volta potential (Eq. 2).
The introduction of plastics, however, can readily overwhelm the Volta potential contribution to flow electrification. Indeed, drops interacting with seemingly identical plastic surfaces gained charges that varied drastically both in polarity and magnitude. The charge across subsequent drops could also grow over time, reminiscent of the explosive growth seen in models of triboelectric charging in granular materials Shinbrot et al. (2017). Additionally, the involvement of non-polar, non-conductive liquids can also result in highly electrified liquid flow. These results suggest that the Volta potential observed in the glass-metal-water system is but one in a myriad of potential electrification mechanisms leading to flow electrification. We hope that future experiments disentangle the particularities of these mechanisms by considering a broader range of solid-fluid (or even fluid-fluid!) pairs. Improved characterizations of these processes will assuredly have important implications for industry, the energy sector, and even planetary science.

### v2

@@ I Introduction
The data for both syringe tips can be fit to a linear relationship. The intercept with the vertical axis ($V_{\text{tip}}=$ 0) agrees well with our measurements using a grounded tip (Fig. 2)–that is, it represents the Volta potential, $\Delta\psi$. Interestingly, the slope of the linear fit is not unity: a small increase in $V_{\text{tip}}$ leads to a slightly smaller increase in $V_{D}$. This result may be expected since the applied voltage can change the distribution of electrons at the surface and the water dipole contribution to the Volta potential in Eq. 2 Mohandas et al. (2024). Although we did not perform independent measurements of $\Delta\psi$ using alternative experimental techniques Le et al. (2017), the consistency of $V_{D}$ over multiple experimental conditions and its quantitative agreement with the expected values of $\Delta\psi$ strongly suggest their equivalency.
In order to obtain a potential difference between the interior bulk water and the ambient air, there must be a net surface charge density at the air-water interface. In water, hydroxide ions (OH-) and hydronium ions (H3O+) are expected to accumulate at the air-water interface Creux et al. (2007). A net negative charge can be due to an imbalance in the natural ion concentration. In pure water, the concentration of each ion species is approximately 1.7 parts per billion (ppb). For a water drop of radius 2 mm, this corresponds to 1.9$\times$1012 ions of each species. Additionally, the charge imbalance required to produce a potential $V_{D}$ = -0.5 V is about 7$\times$105 elementary charges. This means that only a small fraction (less than 1 ppm) of the naturally occurring ions would need to contribute to the net charge in order to produce the expected Volta potential. A potential mechanism to produce this charge imbalance is shown in Fig. 5. A hydronium ion can source an electron from the metal surface, eventually resulting in a minute amount of hydrogen gas that either remains dissolved or escapes into the air. We note that this charge imbalance would produce negligible changes in pH, and thus absorbed gases, such as CO2, could also produce a similar charge imbalance.
@@
Figure 6: PDFs of $\Delta q/\Delta m$ for different fluids dispensed from a 5 cm long, 1.60 or 2.39 mm ID stainless steel luer lock syringe tip. The salt solution and glycerol data were acquired at a flow rate of 350 $\mu$L/min on the 2.39 mm ID tip, while the high density and much lower surface tension of Fluorinert FC-70 required a flow rate of 35 $\mu$L/min on the 1.60 mm ID tip. Each distribution is computed over $n>$ 360 data points. (a) For salt water, $\langle\Delta q\rangle=$0.08 $\pm$ 0.02 pC and $\langle\Delta m\rangle=$ 0.054 $\pm$ 0.004 g. Data were captured at 48% RH. (b) For glycerol, $\langle\Delta q\rangle=$0.13 $\pm$ 0.02 pC and $\langle\Delta m\rangle=$ 0.049 $\pm$ 0.002 g. Data were captured at 48% RH. (c) For Fluorinert FC-70, $\langle\Delta q\rangle=$0.18 $\pm$ 0.02 pC and $\langle\Delta m\rangle=$ 0.016 $\pm$ 0.001 g. Data were captured at 29% RH. Overall, the addition of ions (salt) in the fluid narrows the distribution, whereas non-polar and highly insulating fluids display broader distributions.
@@
The mechanism depicted in Fig. 5 results from the naturally occurring ions and concomitant sub-nanometer Debye length in water. However, the charging mechanisms at play in insulating fluids may be more complex. Thus, using the same glass syringe and stainless steel tips, we also investigated other liquids with the same procedure as for pure water. Figure 6 shows histograms of $\Delta q/\Delta m$ for a 2.5% w/v NaCl aqueous solution, pure glycerol, and a fluorinated hydrocarbon fluid (Fluorinert FC-70). For salt water, both the magnitude and standard deviation of $\Delta q/\Delta m$ were slightly smaller than those of DI water. The presence of free ions in the solution can reduce the potential of zero charge (making it closer to the metal work function), and thus reduces the magnitude of the Volta potential Örnek et al. (2019). Moreover, aqueous ions such as Na+ and Cl- are expected to be less important for interfacial charging mechanisms due to their lack of hydrogen bonding Zimmermann et al. (2001). For glycerol, drop radii ranged from 2–2.2 mm, and $\Delta q/\Delta m$ was close to that of pure water. Although we do not have a prediction for the Volta potential of glycerol, its polar nature may cause it to behave similarly to water. Additionally, glycerol is hygroscopic, suggesting that absorbed water from the ambient environment may cause it to gain charges near the Volta potential for water.
@@
From an applications perspective, the large potentials acquired by fluids with low conductivity evince the persistent electrostatic hazards present across a number of fields. Beyond industrial settings, however, the frictional charging of dielectric fluids may have important implications for geophysical processes on worlds with exotic potamologies. Saturn’s moon Titan, for instance, hosts extensive river and lake systems made not of water, but of liquid methane and ethane Mitri et al. (2007). While previous work has suggested that the charging of hydrocarbon solids can impact the transport of dust on Titan, whether or not flow electrification impacts the transport of hydrocarbon liquids on Titan remains unexplored. Ultimately, we suggest that more experiments are needed to specifically investigate such insulating liquids, whose equilibrium charge may arise from a different balance of charge transfer mechanisms.
Figure 7: Time series of $\Delta q/\Delta m$ for different syringe and tip components using 15 mL of DI water at a flow rate of 350 $\mu$L/min. The expected $\Delta q/\Delta m$ range of -5 pC/g to -1 pC/g for metal and glass is highlighted by the translucent red region on each plot. Vertical axis limits vary between plots to better depict the magnitude of $\Delta q/\Delta m$. (a) Plot of $\Delta q/\Delta m$ versus time for drops dispensed from glass syringe with different metallic syringe tips. Experiments were performed between 30 and 48% RH. The tip length varied, yet all tips had a 2.69 mm ID. Across these experiments, $\langle\Delta m\rangle=$ 0.045 $\pm$ 0.002 g. (b) Plot of $\Delta q/\Delta m$ versus time for drops dispensed from stainless steel tips mounted on PPL syringe bodies. All tips had a length of 5 cm and 1.70 mm ID, and experiments were performed at 46% RH. Across these experiments, $\langle\Delta m\rangle=$ 0.0347 $\pm$ 0.0005 g. (c) Plot of $\Delta q/\Delta m$ versus time for drops dispensed from PTFE tips mounted on a glass syringe body. All tips were cut to a length of 2 cm before Neutrad cleaning and had a 1.35 mm ID, and experiments were performed at $<$ 30% RH. Across these experiments, $\langle\Delta m\rangle=$ 0.0267 $\pm$ 0.0005 g. Overall, the introduction of plastic components reveals time and history dependent charging behavior.
@@
Instruments for handling of liquids (e.g. pipette tips and syringe bodies) commonly employ plastic components. In solid-solid interactions, plastics can often facilitate the transfer of large amounts of electrostatic charge. Beyond our experiments with glass syringes and metal needles, we also investigated the charge gained by water dispensed from a polypropylene (PPL) syringe body with a 5 cm long stainless steel tip, and water deposited from a glass syringe with a PTFE tip. For these experiments, we used syringes sourced from their original sterile packaging without additional cleaning. Figure 7a shows that drops dispensed consecutively from a glass syringe/metal tip combination over a period of $\sim$45 minutes have $\langle\Delta q/\Delta m\rangle=$2.3 $\pm$ 0.3 pC/g. These time series are consistent with the data presented in Fig. 2. In contrast, drops sourced from a PPL syringe and metal tip gained positive charge, with $\Delta q/\Delta m=$ 10–30 pC/g (Fig. 7b). Moreover, we observed an evolution of $\Delta q/\Delta m$ over long timescales, which is solely due to charge variation since the standard deviation of the drop mass was less than 5%. During some trials, $\Delta q/\Delta m$ increased by up to a factor of 2 over the length of the experiment, whereas in other cases we observed a small decrease in $\Delta q/\Delta m$. These experiments show that despite the water passing through 5 cm of metal prior to falling through air, the initial contact with the plastic syringe body dominated the sign, magnitude, and time evolution of the drop charge.
@@
The fact that water drops dispensed from plastic reservoirs or syringe tips gain large amounts of charge is consistent with previous experiments. For example, water drops deposited from a pipette tip can display a large positive charge, $\sim$100 pC or more Choi et al. (2013). A more surprising result is the fact that we observe both negative and positive charging on water droplets interacting with PTFE tips. Like many plastics, PTFE is on the extreme lower end of the triboelectric series and is often considered to be one of the substances that most effectively gains negative charge during frictional interactions. Indeed, the positive electrification of droplets flowing on PTFE surfaces has served as the basis for a number of proposed triboelectric nano generators Kwak et al. (2016); Wu et al. (2021).
Currently, we do not have a satisfactory explanation for the evolution of $\Delta q/\Delta m$ when a plastic component is introduced into the system. The magnitude of the charge we observe in experiments with plastics is 10-100 times larger than with glass and metal (Fig. 2). It is unlikely that an electrochemical charging mechanism similar to Fig. 5 could produce such a large, and often time-varying charge. For solid-solid contacts involving hydrocarbons, microscale chemical heterogeneity along solid surfaces can lead to a large variability in charging behavior Grosjean and Waitukaitis (2023b); Sobolev et al. (2022). Moreover, history dependence in solid-solid tribocharging has recently been demonstrated during repeated contacts between a sphere and a planar surface using acoustic levitation Grosjean and Waitukaitis (2023a). In those experiments, hysteresis in water adsorption was suspected to be the primary cause of history dependence. Surprisingly, materials discharged and retested under the same humidity conditions could produce different magnitudes and signs of charge transfer. Similar variability could operate in solid-liquid contacts, explaining the apparent randomness of charge gain we observed in our experiments with plastics. Lastly, surfactants leached from plastic surfaces could also contribute to the diversity of charging behaviors reported here. We suspect that the matter of water-plastic electrification will find a more satisfactory answer in future experiments considering a broader ensemble of plastics.
@@ Conclusions
By quasistatically depositing individual drops at low flow rates, we show that, for certain material combinations, charging of water can be attributed to well-characterized electrochemical processes. Specifically, the charge on water droplets falling from a glass-metal vessel can be described by the Volta potential (Eq. 2). The introduction of plastics, however, can generate electrification behaviors which deviate substantially from those predicted by the Volta potential. Indeed, drops interacting with seemingly identical plastic surfaces gained charges that varied drastically both in polarity and magnitude. Furthermore, the charge magnitude across subsequent drops could increase over time, reminiscent of the explosive growth seen in models of triboelectric charging in granular materials Shinbrot et al. (2017). Lastly, the involvement of non-polar, non-conductive liquids can also result in highly electrified liquid flow whose behavior warrants targeted studies.
Together, our results suggest that the Volta potential is but one in a myriad of potential electrification mechanisms leading to flow electrification, and motivate future experiments involving a broader range of solid-fluid (or even fluid-fluid) interactions. Furthermore, the varied magnitude and polarity of charging we observe with certain material combinations (e.g., water and plastics) undermines the usefulness of tools like triboelectric series. It has been recently shown that an organized triboelectric series can be built by rubbing identical materials together Sobarzo et al. (2025), implying that history dependence can be much more important for charging behavior than average, bulk material properties. Improved characterizations of fluid electrification phenomena will assuredly resolve these open questions, with implications for industry, the energy sector, and even planetary science. In the meantime, however, our work hints that triboelectric series classification for fluids should be used sparingly and judiciously.